# Training Foundation Models on a Full-Stack AMD Platform: Compute, Networking, and System Design

### v1

@@ Abstract
We report on the first large-scale mixture-of-experts (MoE) pretraining study on pure AMD hardware, utilizing both MI300X GPUs with Pollara interconnect. We distill practical guidance for both systems and model design. On the systems side, we deliver a comprehensive cluster and networking characterization: microbenchmarks for all core collectives (all-reduce, reduce-scatter, all-gather, broadcast) across message sizes and GPU counts on Pollara. To our knowledge, this is the first at this scale. We further provide MI300X microbenchmarks on kernel sizing and memory bandwidth to inform model design. On the modeling side, we introduce and apply MI300X-aware transformer sizing rules for attention and MLP blocks and justify MoE widths that jointly optimize training throughput and inference latency. We describe our training stack in depth, including often-ignored utilities such as fault-tolerance and checkpoint–reshaping, as well as detailed information on our training recipe. We also provide a preview of our model architecture and base model - ZAYA1 (760M active, 8.3B total parameters MoE) - which will be further improved upon in forthcoming papers. ZAYA1-base achieves performance comparable to leading base models such as Qwen3-4B and Gemma3-12B at its scale and larger, and outperforms models including Llama-3-8B and OLMoE across reasoning, mathematics, and coding benchmarks. Together, these results demonstrate that the AMD hardware, network, and software stack are mature and optimized enough for competitive large-scale pretraining.
@@ I Introduction
For this purpose, we introduce ZAYA1-base, a mixture-of-experts (MoE) transformer trained on an MI300X cluster with AMD Pensando Pollara networking. Architecturally, ZAYA1-base follows Zyphra’s novel ‘MoE++’ recipe that combines: (i) Compressed Convolutional Attention (CCA) (cca), which executes attention in a compressed latent space to reduce both prefill compute and KV-cache size; [^1] We use the term prefill to denote the first phase of inference, where the model must first ingest the entire input sequence and generate the first token. This is equivalent to next-token prediction during pretraining. The KV-cache is a purely inference-time structure that stores the KV states of prior tokens so that they do not require recomputation for each subsequent token. (ii) a significantly more powerful and expressive ZAYA1 router that replaces the standard linear gate with a compact MLP with depth-aware mixing to promote expert specialization and stable load balancing; and (iii) lightweight residual scaling that gives the layers fine-grained control over information flow through the residual stream with negligible parameter overhead (See Section IV). Beyond the core model architecture, our training run used context parallelism (CP) (Section V-C1) tailored to CCA, fused-optimizer and layer-norm (LN) kernels (Section V-B), and training/IO paths tuned for the AMD software stack (Sections V-D and V-F). We also optimized the shapes of our model architecture for performance on AMD hardware and describe in detail how to perform this kind of model sizing to optimize training and inference on a particular hardware platform (Sections V-A).
@@ I-A Contributions
  • Networking characterization at scale: We benchmark Pollara/Pensando networking for the collectives that dominate LLM training (e.g., all_reduce, reduce_scatter, all_gather), sweeping message sizes and GPU counts. We relate these measurements to optimizer/gradient communication and to context parallelism (CP) traffic patterns induced by CCA.
@@ II-A Node Architecture
Each compute node consists of eight MI300X GPUs connected via InfinityFabric, dual-socket Intel Xeon CPUs with 2 TB of DDR5 memory, a dedicated Pollara 400 Gbps NIC per GPU, and local NVMe storage for high-throughput datasets and checkpoints (see Figure 2). A detailed description for all node types is provided in Table I and Appendix A.
@@ II-B Cluster Architecture
Our cluster employs two physically separate networks. The Pollara training fabric handles all collective communication for distributed training, while a VPC network manages dataset I/O, checkpoints, and cluster management. This separation prevents storage operations from interfering with model communication (gradients, optimizer states, etc). The VPC network connects each node via Pensando DSC SmartNIC/DPUs to a separate leaf–spine fabric, providing approximately 200 Gbps throughput to the storage node.
@@ III-A GPU HBM Bandwidth
Choosing a fair software setup and benchmarking regime for measuring HBM bandwidth, as with any other hardware metric, is challenging. In order to test the hardware’s limits, vendors and kernel engineers opt for low-level benchmarks with custom data transfer kernels such as (nvbandwidth; rocm_systems). When conveying application results such as DL frameworks, these low-level workloads are often not representative of the available memory transfer kernels, even when under full load. However, if application benchmarks are exclusively relied upon, one is indirectly measuring both: 1) How well that application is optimized for the hardware in question, which heavily favors the incumbent, and 2) What overhead the library incurs. Note that the severity of these effects differs per hardware element and its underlying kernels. Specifically, while PyTorch communication operations and compute GEMMs are nearly a direct mapping to their underlying libraries (e.g. RCCL and rocBLAS, respectively), hardware like HBM and storage (NVMe/SSDs) must be handled more carefully.
With this in mind, we implemented a PyTorch benchmark that mirrors the memory access patterns of a benchmark such as babelstream by carefully laying out the tensor when reading. Further, we tuned the memcpy kernel within PyTorch to better select the kernel thread block size depending on the tensor size for the MI300X backend. While not all of the memory access patterns incurred by our workloads are so clean (e.g. non-contiguous tensors and sparse access patterns), we believe these results to better capture what the HBM hardware is capable of under real training and inference workloads. The results of which are depicted in Figure 3.
Figure 3: The achievable memory bandwidth to HBM for PyTorch using the ROCm backend.
@@ III-B GPU Compute
Most of a model’s iteration time is spent performing BF16 GEMM kernels (see Figure 10). However, not all GEMM shapes are created equal—some shapes from a given model are far more amenable to a GPU’s hardware than others (see Table III and (anthony2024codesign)). This makes sizing the model appropriately, so that the underlying GEMM shapes are performant on the target inference hardware, an extremely important step in model design. To optimize GEMM performance for ZAYA1, we adopted a systematic approach. We began by performing an exhaustive search over all possible M, N, and K matrix shapes and their associated GEMM libraries (see Figure 4). Beyond shape sizing, we also addressed the selection of GEMM backends and algorithms. ROCm provides multiple GEMM backends (rocBLAS and hipBLASLt), each containing many algorithms. We performed static tuning via a combination of PyTorch TunableOp, ROCm TransformerEngine (transformer_engine), and HIPBLASLt-bench tune from Primus primus2025. This tuning produced static lookup tables that map GEMM sizes to the most performant algorithms within rocBLAS and hipBLASlt, which are then loaded at runtime to ensure optimal algorithm selection. We then conducted a series of sizing sweeps across each component block in our model, allowing us to round the hyperparameters chosen for ZAYA1 (see Table II) to efficient sizes. Our findings confirmed that larger GEMMs perform better—on the MI300X, a problem size of approximately 200 GFLOPs is required to reach peak throughput.
@@ III-C InfinityFabric Bandwidth
Since the growth in compute per GPU has far outpaced inter-node interconnect bandwidth per node, DL clusters require intra-node high-bandwidth interconnects such as AMD InfinityFabric (amd2024cdna3) or NVIDIA NVLINK (nvidia_nvlink). These high-bandwidth intra-node interconnects are necessary to support parallelism schemes that require high communication volume such as expert- and tensor-parallelism (gshard; megatron; anthony2024comms). This allows the inter-node interconnect to be used for less bandwidth-intensive communication operations such as data parallelism or pipeline parallelism (rajbhandari2020zero; gpipe). Additionally, high-bandwidth intra-node interconnects enable collective communication libraries (e.g. RCCL/NCCL/MPI) to design two-level collective algorithms that rely heavily on the intra-node fabric, thus indirectly enabling scale-out performance on the inter-node fabric.
Most DL software makes the assumption of a switched intra-node topology like NVSwitch. This topology allows for full bandwidth for any world size in a given communication operation. However, AMD’s InfinityFabric uses xGMI (xgmi), which requires all GPUs participate in a given collective operation to achieve full bandwidth (see Eq. 1). Parallelism schemes and DL implementations for current AMD clusters must therefore account for this. Our initial approach is that parallelism degrees should either span a full node or not be used at all. By enforcing this requirement, we can reformulate classically point-to-point schemes as collective operations. An example of this is tree attention (shyam2024tree), which reformulates the attention operation such that the point-to-point message passing approach of ring-attention (liu2023ring) can be replaced by all-reduce, an all-to-all operation.
@@
Where $n$ is the number of GPUs participating in the intra-node communication operation ($1\leq n\leq 8$), $B_{max}$ is the maximum achievable bandwidth ($\approx 450$ GBps for MI300X xGMI), and $B_{link}$ is the bandwidth of each xGMI link ($64$ GBps for our MI300X xGMI fabric).
We depict results for intra-node communication operations in Figures 5. Our parallelism topology for ZAYA1 is just the ZeRO-1 distributed optimizer during initial pretraining at 4096 sequence length, and ZeRO-1 + context parallelism (ring attention (liu2023ring)) when scaling context up to 32768. While the All_Gather and All_Reduce (ZeRO-1) operations were two-level algorithms, we incurred many point-to-point operations within the node when performing context parallelism during pretraining. Long-context inference relies heavily on tree attention (shyam2024tree) to avoid the bandwidth bottleneck of xGMI. This is because backwards compute will not be available for additional overlap, and forwards compute is not sufficient for full point-to-point overlap on xGMI.
@@ III-D Pollara Bandwidth
However, interconnects and their associated communication libraries require large messages to fully utilize the interconnect’s bandwidth. Small messages are latency-bound (meaning that they spend the majority of their time in fixed per-message overheads such as kernel launch, synchronization, switch traversal, etc). Larger messages are bandwidth-bound (meaning that the one-time costs are amortized, and more bandwidth-efficient collective algorithms such as message pipelining may be used). See Figure 7 for examples of this effect.
@@ IV-B Training
Following recent works (gu2025jet; bakouch2025smollm3), ZAYA1-base was trained in three phases. During phase 1, which lasted for 8T tokens, ZAYA1-base was pretrained on a mix that consisted primarily of web-crawl data with additional code, math, and multilingual data mixed in. This was followed by a 4T-token phase 2 in which the proportion of code and math data was significantly increased. Further, additional reasoning and SFT-style data was added, with reasoning traces trimmed to preserve full answers within the 4k context window. During these two phases, the learning rate schedule was a cosine decay from 6e-4 to 2e-4. During the mid-training phase 3, we used a cosine decay from 2e-4 to 1.5e-4 during context extension. This was followed by our final annealing phase, for which we used a linear decay to 7e-6 to prepare for post-training. See Figure 9 for our full training schedule to produce ZAYA1-base.
@@
During phase 1 of pretraining, we were able to fit a minibatch size of 5 in HBM. During context extension to 32k, we trained with minibatch size 1 and with context parallel size of 2. For 128k context, we merely require training with a context-parallel size of 8 and do not yet require inter-node communication. The substantial reductions in attention FLOPs and activation memory that CCA provided also made context extension significantly easier than for alternative attention methods since both our training FLOPs and activation memory were reduced by approximately $8\times$ compared to e.g. MLA. This allowed us to train at 32k context with approximately the same efficiency at 4k context.
@@ V-A Model Sizing
We have broken down the iteration time during core ZAYA1 pretraining (sequence length 4096), and depict the results in Figure 10. Breakdowns are collected by summing and tagging GPU operations from a PyTorch profile. ”Optimizer comm” and ”Optimizer compute” denote the communication of our distributed Muon implementation and its inherent Newton-Shulz (NS) iterations. While forward and backward computation in the core kernels (i.e. convolutions and softmax attention for CCA, expert MLP GEMMs, etc) takes the majority of iteration time, computation and communication overhead from the distributed Muon optimizer contribute a non-negligible effect. This optimizer compute overhead motivated the Muon kernel in Section V-B1, and the communicated overhead motivated the tensor fusion tuning described in Sections III-D and III-C.
@@ V-B Kernels
We wrote custom HIP kernels for the layer-norm operation and for several subroutines used in the Muon optimizer.
@@ V-C Parallelism
V-C1 Context Parallelism
In order to extend the context of ZAYA1 base, we use context parallelism to shard the sequence among multiple GPUs. Specifically, we shard the sequence into $2\times(\textrm{context parallel world size})$ GPUs where each GPU holds $(\textrm{sequence length})/(2\times(\textrm{context-parallel world size}))$ tokens. These shards are shuffled in order to maintain load balancing during the causal attention computation. We use the context parallel attention algorithm known as Ring Attention liu2023ring which sends key and value shards in a point to point manner between context parallel ranks that form a logical ring pattern. Local attention is computed and renormalized as each rank receives new key and value shards. This computation is overlapped with communication.

### v2

@@ Abstract
We report on the first large-scale mixture-of-experts (MoE) pretraining study on pure AMD hardware, utilizing both MI300X GPUs and Pollara networking. We distill practical guidance for both systems and model design. On the systems side, we deliver a comprehensive cluster and networking characterization: microbenchmarks for all core collectives (AllReduce, ReduceScatter, AllGather, Broadcast) across message sizes and GPU counts over Pollara. To our knowledge, this is the first at this scale. We further provide MI300X microbenchmarks on kernel sizing and memory bandwidth to inform model design. On the modeling side, we introduce and apply MI300X-aware transformer sizing rules for attention and MLP blocks and justify MoE widths that jointly optimize training throughput and inference latency. We describe our training stack in depth, including often-ignored utilities such as fault-tolerance and checkpoint–reshaping, as well as detailed information on our training recipe. We also provide a preview of our model architecture and base model - ZAYA1 (760M active, 8.3B total parameters MoE, available at https://huggingface.co/Zyphra/ZAYA1-base) - which will be further improved upon in forthcoming papers. ZAYA1-base achieves performance comparable to leading base models such as Qwen3-4B and Gemma3-12B at its scale and larger, and outperforms models including Llama-3-8B and OLMoE across reasoning, mathematics, and coding benchmarks. Together, these results demonstrate that the AMD hardware, network, and software stack are mature and optimized enough for competitive large-scale pretraining.
@@ I Introduction
For this purpose, we introduce ZAYA1-base (https://huggingface.co/Zyphra/ZAYA1-base), a mixture-of-experts (MoE) transformer trained on an MI300X cluster with AMD Pensando Pollara networking. Architecturally, ZAYA1-base follows Zyphra’s novel ‘MoE++’ recipe that combines: (i) Compressed Convolutional Attention (CCA) (cca), which executes attention in a compressed latent space to reduce both prefill compute and KV-cache size; [^1] We use the term prefill to denote the first phase of inference, where the model must first ingest the entire input sequence and generate the first token. This is equivalent to next-token prediction during pretraining. The KV-cache is a purely inference-time structure that stores the KV states of prior tokens so that they do not require recomputation for each subsequent token. (ii) a significantly more powerful and expressive ZAYA1 router that replaces the standard linear gate with a compact MLP with depth-aware mixing to promote expert specialization and stable load balancing; and (iii) lightweight residual scaling that gives the layers fine-grained control over information flow through the residual stream with negligible parameter overhead (See Section IV). Beyond the core model architecture, our training run used context-parallelism (CP) (Section V-C1) tailored to CCA, fused-optimizer and LayerNorm (LN) kernels (Section V-B), and training/IO paths tuned for the AMD software stack (Sections V-D and V-F). We also optimized the shapes of our model architecture for performance on AMD hardware and describe in detail how to perform this kind of model sizing to optimize training and inference on a particular hardware platform (Sections V-A).
@@ I-A Contributions
  • Networking characterization at scale: We benchmark Pollara/Pensando networking for the collectives that dominate LLM training (e.g., AllReduce, ReduceScatter, AllGather), sweeping message sizes and GPU counts. We relate these measurements to optimizer/gradient communication and to context-parallelism (CP) traffic patterns induced by CCA.
@@ II-A Node Architecture
Each compute node consists of eight MI300X GPUs connected via InfinityFabric, dual-socket Intel Xeon CPUs with 2 TB of DDR5 memory, a dedicated Pollara 400 Gbps NIC per GPU, and local NVMe storage for high-throughput datasets and checkpoints (see Figure 2). A detailed description for the compute, storage, and login nodes is provided in Table I and Appendix A.
@@ II-B Cluster Architecture
Our cluster employs two physically separate networks. The Pollara training fabric handles all collective communication for distributed training, while a VPC network manages dataset I/O, checkpoints, and cluster management. This separation prevents storage operations from interfering with model communication (gradients, optimizer states, etc). The VPC network connects each node via Pensando DSC SmartNIC/DPUs to a separate leaf–spine fabric, providing approximately 200 Gbps throughput to the storage node from any of the compute nodes.
@@ III-A GPU HBM Bandwidth
Choosing a fair software setup and benchmarking regime for measuring HBM bandwidth, as with any other hardware metric, is challenging. In order to test the hardware’s limits, vendors and kernel engineers opt for low-level benchmarks with custom data transfer kernels such as nvbandwidth (nvbandwidth) and the ROCm bandwidth tests (RBT) (rocm_systems). When conveying application results such as DL frameworks, these low-level workloads are often not representative of the available memory transfer kernels, even when under full load. However, if application benchmarks are exclusively relied upon, one is indirectly measuring both: 1) How well that application is optimized for the hardware in question, which heavily favors the incumbent, and 2) What overhead the library incurs. Note that the severity of these effects differs per hardware element and its underlying kernels. Specifically, while PyTorch communication operations and compute GEMMs are nearly a direct mapping to their underlying libraries (e.g. RCCL and rocBLAS, respectively), hardware like HBM and storage (NVMe/SSDs) must be handled more carefully.
With this in mind, we implemented a PyTorch benchmark that mirrors the memory access patterns of a benchmark such as (babelstream) by carefully laying out the tensor when reading. Further, we tuned the memcpy kernel within PyTorch to better select the kernel thread block size depending on the tensor size for the MI300X backend. While not all of the memory access patterns incurred by our workloads are so clean (e.g. non-contiguous tensors and sparse access patterns), we believe these results to better capture what the HBM hardware is capable of under real training and inference workloads. The results of which are depicted in Figure 3.
Figure 3: The achievable memory bandwidth to HBM for PyTorch using the ROCm/CUDA backends.
@@ III-B GPU Compute
Most of a model’s iteration time is spent performing BF16 GEMM11footnotetext: GEMM (General Matrix Multiplication) computes $C=\alpha AB+\beta C$, where $A$ is an $M\times K$ matrix, $B$ is a $K\times N$ matrix, and $C$ is an $M\times N$ matrix, with $\alpha$ and $\beta$ as scalar coefficients. The operation requires $2MNK$ FLOPs (floating-point operations), as each of the $MN$ output elements involves a dot product of $K$ elements. kernels (see Figure 10). However, not all GEMM shapes are created equal—some shapes from a given model are far more amenable to a GPU’s hardware than others (see Table III and (anthony2024codesign)). This makes sizing the model appropriately, so that the underlying GEMM shapes are performant on the target inference hardware, an extremely important step in model design. To optimize GEMM performance for ZAYA1, we adopted a systematic approach. We began by performing an exhaustive search over all possible M, N, and K matrix shapes and their associated GEMM libraries (see Figure 4). Beyond shape sizing, we also addressed the selection of GEMM backends and algorithms. ROCm provides multiple GEMM backends (rocBLAS and hipBLASLt), each containing many algorithms. We performed static tuning via a combination of PyTorch TunableOp, ROCm TransformerEngine (transformer_engine), and HIPBLASLt-bench tune from Primus (primus2025). This tuning produced static lookup tables that map GEMM sizes to the most performant algorithms within rocBLAS and hipBLASlt, which are then loaded at runtime to ensure optimal algorithm selection. We then conducted a series of sizing sweeps across each component block in our model, allowing us to round the hyperparameters chosen for ZAYA1 (see Table II) to efficient sizes. Our findings confirmed that larger GEMMs perform better—on the MI300X, a problem size of approximately 200 GFLOPs is required to reach peak throughput.
@@ III-C InfinityFabric Bandwidth
Since the growth in compute per GPU has far outpaced inter-node interconnect bandwidth per node, DL clusters require intra-node high-bandwidth interconnects such as AMD InfinityFabric (amd2024cdna3) or NVIDIA NVLINK (nvidia_nvlink). These high-bandwidth intra-node interconnects are necessary to support parallelism schemes that require high communication volume such as expert- and tensor-parallelism (gshard; megatron; anthony2024comms). This allows the inter-node interconnect to be used for less bandwidth-intensive communication operations such as data parallelism or pipeline-parallelism (rajbhandari2020zero; gpipe). Additionally, high-bandwidth intra-node interconnects enable collective communication libraries (e.g. RCCL/NCCL/MPI) to design two-level collective algorithms that rely heavily on the intra-node fabric, thus indirectly enabling scale-out performance on the inter-node fabric.
Most DL software makes the assumption of a switched intra-node topology like NVSwitch. This topology allows for full bandwidth for any world size in a given communication operation. However, AMD’s InfinityFabric uses xGMI (xgmi), which requires all GPUs participate in a given collective operation to achieve full bandwidth (see Eq. 1). Parallelism schemes and DL implementations for current AMD clusters must therefore account for this. Our initial approach is that parallelism degrees should either span a full node or not be used at all. By enforcing this requirement, we can reformulate classically point-to-point schemes as collective operations. An example of this is tree attention (shyam2024tree), which reformulates the attention operation such that the point-to-point message passing approach of ring-attention (liu2023ring) can be replaced by AllReduce, an All-to-All operation.
@@
Where $n$ is the number of GPUs participating in the intra-node communication operation ($1\leq n\leq 8$), $B_{\text{max}}$ is the maximum achievable bandwidth ($\approx 450$ GBps for 8-MI300X communication across all xGMI links), and $B_{link}$ is the bandwidth of each xGMI link ($64$ GBps for our MI300X xGMI fabric).
We depict results for intra-node communication operations in Figures 5. Our parallelism topology for ZAYA1 is just the ZeRO-1 distributed optimizer during initial pretraining at 4096 sequence length, and ZeRO-1 + context-parallelism (ring attention (liu2023ring)) when scaling context up to 32768. While the AllGather and AllReduce (ZeRO-1) operations were two-level algorithms, we incurred many point-to-point operations within the node when performing context-parallelism during pretraining. Long-context inference relies heavily on tree attention (shyam2024tree) to avoid the bandwidth bottleneck of xGMI. This is because backwards compute will not be available for additional overlap, and forwards compute is not sufficient for full point-to-point overlap on xGMI.
@@ III-D Pollara Bandwidth
However, interconnects and their associated communication libraries require large messages to fully utilize the interconnect’s bandwidth. Small messages are latency-bound (meaning that they spend the majority of their time in fixed per-message overheads such as kernel launch, synchronization, switch traversal, etc). Larger messages are bandwidth-bound (meaning that the one-time costs are amortized, and more bandwidth-efficient communication algorithms such as message pipelining may be used). See Figures 6 and 7 for examples of this effect.
@@ IV-B Training
Following recent works (gu2025jet; bakouch2025smollm3), ZAYA1-base was trained in three phases. During phase 1, which lasted for 8T tokens, ZAYA1-base was pretrained on a mix that consisted primarily of web-crawl data with additional code, math, and multilingual data mixed in. This was followed by a 4T-token phase 2 in which the proportion of code and math data was significantly increased. Further, additional reasoning and SFT-formatted data was added, with reasoning traces trimmed to preserve full answers within the 4k context window. During these two phases, the learning rate schedule was a cosine decay from 6e-4 to 2e-4. During the mid-training phase 3, we used a cosine decay from 2e-4 to 1.5e-4 during context extension. This was followed by our final annealing phase, for which we used a linear decay to 7e-6 to prepare for post-training. See Figure 9 for our full training schedule to produce ZAYA1-base.
@@
During phase 1 of pretraining, we were able to fit a minibatch size of 5 in HBM. During context extension to 32k, we trained with minibatch size 1 and with context-parallel size of 2. For 128k context, we merely require training with a context-parallel size of 8 and do not yet require inter-node communication. The substantial reductions in attention FLOPs and activation memory that CCA provided also made context extension significantly easier than for alternative attention methods since both our training FLOPs and activation memory were reduced by approximately $8\times$ compared to e.g. MLA. This allowed us to train at 32k context with approximately the same efficiency at 4k context.
@@ V-A Model Sizing
We have broken down the iteration time during core ZAYA1 pretraining (sequence length 4096), and depict the results in Figure 10. Breakdowns are collected by summing and tagging GPU operations from a PyTorch profile. ”Optimizer comm” and ”Optimizer compute” denote the communication of our distributed Muon implementation and its inherent Newton-Schulz (NS) iterations. While forward and backward computation in the core kernels (i.e. convolutions and softmax attention for CCA, expert MLP GEMMs, etc) takes the majority of iteration time, computation and communication overhead from the distributed Muon optimizer contribute a non-negligible effect. This optimizer compute overhead motivated the Muon kernel in Section V-B1, and the communicated overhead motivated the tensor fusion tuning described in Sections III-D and III-C.
@@ V-B Kernels
We wrote custom HIP kernels for the LayerNorm operation and for several subroutines used in the Muon optimizer.
@@ V-C Parallelism
V-C1 Context-Parallelism
In order to extend the context of ZAYA1 base, we use context-parallelism to shard the sequence among multiple GPUs. Specifically, we shard the sequence into $2\times(\textrm{context-parallel world size})$ GPUs where each GPU holds $(\textrm{sequence length})/(2\times(\textrm{context-parallel world size}))$ tokens. These shards are shuffled in order to maintain load balancing during the causal attention computation. We use the context-parallel attention algorithm known as Ring Attention (liu2023ring) which sends key and value shards in a point to point manner between context-parallel ranks that form a logical ring pattern. Local attention is computed and renormalized as each rank receives new key and value shards. This computation is overlapped with communication.